what is the slope of a line that is parallel to the line represented by the equation 4y-5x=6
Answer by J2R2R(94)

What is the slope of a line that is parallel to the line represented by the equation 4y – 5x = 6?

4y – 5x = 6 can be written as

4y = 5x + 6 or y = 5x/4 + 3/2

For the line y = mx + c the slope or gradient is m, so for our equation we have a slope of 5/4.

Any line parallel to this will also have the same slope.

Therefore the slope of the line that is parallel to the line represented by the equation 4y – 5x = 6 is 5/4.